A project has the following cash flows. Co C1 C2 C3 ($700) $200 $500 $244 a. What is the projects payback period? Round the answer to two decimal places. 2 V year(s) b. Calculate the projects NPV at 9%. Do not round intermediate calculations. Round PVF and PVFA values in intermediate calculations to four decimal places. Round the answer to two decimal places C. Calculate the projects PI at 9%. Do not round intermediate calculations. Round PVF and PVFA values in intermediate calculations to four decimal places. Round the answer to two decimal places.
Answer #1

a)

Cumulative cash flow for year 0 = -700

Cumulative cash flow for year 1 = -700 + 200 = -500

Cumuative cash flow for year 2 = -500 + 500 = 0

Project's payback is 2 years

b)

NPV = Present value of cash inflows - present value of cash outflows

NPV = -700 + 200 / ( 1 + 0.09)1 + 500 / ( 1 + 0.09)2 + 244 / ( 1 + 0.09)3

NPV = -700 + 792.739

NPV = $92.74

c)

PI = Present value of cash inflows / initial investment

PI = 792.739 / 700

PI = 1.13